what is the lenght of the conjugate axis?
(x-1)^2/25 -(y+3)^2/9=1
1
Expert's answer
2020-11-15T18:04:59-0500

"\frac{(x-1)^2}{25}- \frac{(y+3)^2}{9}=1" ........ (1)

The standard equation of the hyperbola is given by:

"\frac{(x-h)^2}{a^2}-\frac{(y-k)^2)}{b^2}=1"

When we compare the standard equations and equation (1) we obtain

"a^2= 25" and "b^2=9"

This implies that "a=5" and "b=3"

Conjugate axis"=2b"

"=2 \times 3=6" units